OpenStudy (anonymous):

the perimeter of a rectangle is 176 inches the length exceeds the width by 8 inches find the length and width answer must be a decimal

OpenStudy (anonymous):

\[length = width + 8\]\[perimeter = 2*length + 2*width\]\[perimeter = 2 * (width + 8) + 2* width\]\[perimeter = 4*width + 16 = 176\]\[4*width = 160\]\[width = 40\]\[length = 48\]

OpenStudy (anonymous):

176=(L+W)2 176/2=L+W L+W=88 L= 8+W 8+W+W=88 2W=88-8 W=80/2 W=40 L+40=88 L=88-40 L=48